OpenStudy (anonymous):

Is 0=50 a Contradiction, Conditional equation or Identity?

OpenStudy (mertsj):

Does 0 =50?

OpenStudy (anonymous):

well the actual problem is -4(x - 5) - 61 = 5x - 9(x - 1)

OpenStudy (mertsj):

\[-4x+20-61=5x-9x+9\] \[-41=9\]

OpenStudy (mertsj):

That is never going to be true no matter what value you use for x so I guess that is a contradiction.